# Brain Organoid Computing – an Overview

### v1

@@ II-B4 Interfacing
Current brain organoid computing experiments mainly rely on observing changes in the pattern of elicited spikes [jordan_open_2024]. Typically, predefined regions of activity are read out and mapped according to predetermined rule sets [bakkum_spatio-temporal_2008, kagan_vitro_2022]. Thus, the readout resolution is still quite limited and requires further research to make better use of brain organoids and their capabilities. The same holds true for applying input data to an organoid.
Most current experiments rely on reinforcing organoid activity with external electrical stimuli emulating rewards or punishments to train them to produce desired neural activity [bakkum_spatio-temporal_2008]. However, there is no general method of reliably performing input-output operations with brain organoids [jordan_open_2024].
Current reinforcement training involves providing the brain organoids with predictable stimuli [^2] In this case 100Hz for 100ms. , typically with short, specific high-frequency signals, to targeted cell regions for positive feedback, and unpredictable stimuli [^3] In this case 5Hz for 4s., characterized by ”random” regions being stimulated with longer, low-frequency signals, for negative feedback[kagan_vitro_2022].
In addition to this, random spikes elicited by neurons generate noise, which can result in unreliable and variable results [ruaro_toward_2005], further contributing to the issue of interpreting neural activity and interfacing with the organoids. There is early interfacing software, which aims to solve these issues and provide the capabilities of performing described experiments. There are early softwares that aim to simplify interfacing with brain organoids, for example enabling the brain organoids to control external agents [tessadori_closed-loop_2015] and to be addressed remotely over the internet [jordan_open_2024].
@@ II-B5 Ethics
The foremost ethical concern regarding experimentation on human brain organoids is the possibility of them being conscious. Teams of researchers have conducted analyses of experiments and research data to try to determine if current brain organoids display signs of higher-order thinking. Current research states that these organoids do not possess the complexity to allow for higher-order brain function, associated with consciousness [boyd_moral_2024]. Researchers also argue that the ability to learn and complex activity signatures by themselves is not a sign of consciousness and that without a corporeal state, consciousness cannot arise [croxford_case_2024]. However, experiments showcasing brain organoids and their plasticity, characterize them as posessing a basic form of sentience [kagan_vitro_2022]. This does not constitute consciousness, as the concept remains undefined with several competing theoretical models. Currently, the only widely acknowledged consciousness theory that allows such brain organoids to be classified as being some level of conscious, is the integrated information theory (IIT) [montoya_what_2023]. It states that any system capable of integrating information and learning independently, is at least somewhat conscious [sattin_theoretical_2021].
The experiments mentioned before demonstrate this learning capability of brain organoids [kagan_vitro_2022, bakkum_spatio-temporal_2008, tessadori_closed-loop_2015, cai_brain_2023]. Following the IIT, brain organoids qualify as being at least somewhat conscious [montoya_what_2023]. However, this assessment itself is contested as some researchers in the field argue that the IIT and its representation is basically flawed for accurately depicting possible states of consciousness in in vitro cultures [kagan_neurons_2022]. The definition of consciousness remains open, with numerous contradictory models and theories [sattin_theoretical_2021, croxford_case_2024], leaving the question unresolved whether brain organoids can feel pain or process complex emotions and sensations.
Due to ethical concerns, some countries, such as the U.S.A. and Germany, impose regulations on research with human brain organoids, which could limit research in the future. However, in said countries, no laws prohibit research with human brain organoids. Nonetheless, guidelines for experimentation with human stem cells, must be followed and monitored [pichl_ethical_2023]. If consciousness is indeed discovered in more complex brain organoids, stricter regulations could slow or even halt progress and future organoid-based computers might be entirely prohibited.
@@ II-B6 Complexity
Though brain organoids exhibit complex neural structures, they are obviously by far not as complex as fully developed human brains and can be compared with early stages of embryonic development [qian_brain_2019] at best. This limits their current processing potential. Thus, contemporary brain organoids are vastly inferior in terms of processing power to traditional processing methods and will most likely continue to be until more developed organoids can be grown.
In addition to this writing, reading, and interpreting information to and from such organoids remains an additional challenge that hinders current developments [cai_brain_2023]. While certain promising capabilities of BOs have been demonstrated experimentally, the full integration and understanding of these systems remain major challenges. Thus, a better understanding of neural structures, their processing of stimuli, and the signals they generate is necessary. Current challenges of brain organoids for use in computing culminate in one word: complexity. The technology comes with various caveats, large challenges all by themselves, resulting in an incredibly complex set of problems to solve, if researchers want to fully understand and make use of brain organoids in a computing context, being the major stumbling block for current applications of brain organoids in a computational context.
@@ III-A Artificial Intelligence
Biological neural networks like human brains can be seen as larger, more complex and ANNs with an extremely high degree of interconnectivity [stoop_note_2021], which formed the conceptual basis for digital systems such as ChatGPT, Google Gemini and others. The use of brain organoids, too, could heavily influence the field of AI in the next coupled of years. In fact, human brain organoids have been used to conduct traditionally AI related tasks.
@@
Basic image processing [ruaro_toward_2005] as well as the prediction of the behavior of dynamic systems exhibiting chaotic behavior [cai_brain_2023] are typical tasks that to which in vitro neurons have been applied.
@@
For the latter, such brain organoids required only 4 training epochs to achieve comparable predictive performance, whereas a machine learning algorithm with long short-term memory (LSTM) needed 50 epochs to outperform the biological system, representing a $>$90% training time decrease [cai_brain_2023].
@@
For the latter, such organoids were able to be trained in a much shorter period of time, The demonstrated ability of human brain organoids to adapt to stimuli makes these functionally capable of solving more complex tasks typically associated with machine learning powered AI, such as speech recognition [cai_brain_2023] or actively controlling systems such as video games [kagan_vitro_2022].
@@
The inherent capability of organoids to change their connection structure makes them – at least in perspective – more powerful than traditional ANNs. Contemporary ANNs being larger and computationally intensive, they require pruning and specialization for efficient deployment [he_structured_2023]. In contrast to this, the self-organized, redundant, and interconnected nature of brain organoids lends itself to being better suited for generalist AI systems [stoop_note_2021, jeon_distinctive_2023]. Human brain organoids could therefore be trained in shorter periods, given some major advances in interfacing technology, provide efficient AI solutions and complete tasks in parallel due to the neuronal structure that constitutes them. Their ability to learn autonomously opens possibilities for the use of this technology for a wide range of tasks that currently rely on machine learning, with BNNs potentially surpassing the abilities of contemporary, artificial systems.
@@
Another area of application is the integrations of classic ANNs with such organoids in order to complement each other’s strengths and weaknesses. This has been coined as organoid intelligence (OI) [smirnova_organoid_2023, basu_organoid_2024].
@@
Analyzing vast amounts of data remains a strong suit of ANNs [badai_review_2020] and can be leveraged to characterize and analyze the vast amounts of data generated by brain organoids. This might be used in turn to better understand BNNs and their characteristics. Supplementing ANNs that excel at specialized tasks [jeon_distinctive_2023] with more generalized intelligences through the application of brain organoids might eventually lead to significant advances in the field of AI [cai_brain_2023, basu_organoid_2024]. Being able to learn from minimal sets of data, combined with the analytical strength of contemporary AI systems can enable superior predictive models, higher level decision-making and data analysis models [basu_organoid_2024] (organoid intelligence).
@@
Potential real world applications have been proposed and demonstrated in the field of robotics [polykretis_astrocyte-modulated_2020, independent_organoid_robot, hartung_organoid_2024], AI algorithms in general [cai_brain_2023, smirnova_organoid_2023], such as those used for natural language processing, etc.
@@ III-B Neuromorphic Computing Research
Current neuromorphic processing units are limited by the available processing power with respect to the scope of the brain regions that they can currently simulate. Only partial (tiny) simulations of brain functionality [cai_brain_2023] are possible today. Though brain organoids similarly do not represent fully developed brains and a fully diverse set of cells, current brain organoids are still more representative of portraying true brain functionality than any artificial means developed as of yet [qian_brain_2019]. Accordingly, they can be used as inspiration for improved neuromorphic systems. The current and potential benefits of brain organoids can be leveraged for progress on more efficient, adaptable, and biologically representative neuromorphic processing units, see [thakur_large-scale_2018].
Contemporary research efforts in the area of neuromorphic processors almost exclusively stick to digital computer architectures, thus limiting the plasticity of the neuromorphic architectures under consideration [schuman_opportunities_2022]. However, there are several recent efforts to provide such systems with new substrates that can better mimic the plasticity of BNNs, with the use memristors being one of the more widespread approaches [zhong_dynamic_2021, schuman_opportunities_2022]. The ability of memristive devices to change their conductance according to basically the time integral over the current flowing through the device (there are devices exhibiting thresholds to distinguish between writing and reading) and store this state [zhong_dynamic_2021], makes them highly interesting for research on neuromorphic systems. Memristors are currently used in tasks requiring autonomous learning and it has been demonstrated that on-chip/on-the-edge learning is a viable area of application. One particular example showed memristor based integrated circuits which learned to control a remote controlled car with input provided through a camera [zhang_edge_2023]. Other examples include the recognition of written and spoken digits or words as well as the solution of dynamic systems [hu_memristorbased_2018, zhong_dynamic_2021].
@@
Another key characteristic of biological systems is that of cell diversity. Contemporary neuromorphic architectures focus mainly on emulating neurons in a very simplified way, neglecting the computational functions provided by other cell types critical to cognition [schuman_opportunities_2022, cai_brain_2023]. For instance the actual behavior of dendrites is crucial in the processing of information within biological neuron, however, they are modeled in a simplified way, not fully emulating their actual function [acharya_dendritic_2022]. They have been shown to be paramount in enabling continued learning and long-term plasticity [acharya_dendritic_2022]. Employing better models could possibly alleviate the issue of catastrophic forgetting in ANNs, a problem for which currently no optimal solution exists [jedlicka_contributions_2022].
In addition to this, quite a lot of different cell types are completely missing from current neuromorphic systems, such as astrocytes, which play important roles in network plasticity and regulating the spiking behavior of neurons [poli_experimental_2019]. Some experiments aim to simulate and integrate such missing cell types, though research in this direction is still in its infancy [poli_experimental_2019]. The plasticity of BNNs stems – at least partially – from these different cells and the inherent prediction of possible future stimuli [luczak_neurons_2022, basu_organoid_2024]. In contrast to these contemporary digital approaches, brain organoids already contain these cells, making them more biologically representative than contemporary neuromorphic systems [qian_brain_2019, kagan_vitro_2022, jordan_open_2024].
Given the limitations of current neuromorphic systems in replicating biological complexities, their adaptability to uncertainty and real-world scenarios is significantly hampered. This is partly due to the absence of a comprehensive biological representation of neural structures [cai_brain_2023]. Furthermore, while spiking associated with BNNs enhances their efficiency [schuman_opportunities_2022, sharf_functional_2022], most of neuromorphic systems of today are not of the spiking type [hu_memristorbased_2018, zhong_dynamic_2021, zhang_edge_2023], limiting their potential efficiency, although there are exceptions [poli_experimental_2019, schuman_opportunities_2022].
@@ IV Key Research
Maybe the most important area of research is the question of how to provide brain organoids with more abstract information and how to extract and interpret their responses. Solving these problems will advance their application in advanced AI use cases in general [jordan_open_2024]. Consequently, research must focus on advancing MEA technology. This is crucial to more finely discern spikes from the organoids, in addition to allowing more precise addressing of their different cell structures. Current MEAs are severely limited with respect the amount of addressable cells [cai_brain_2023]. One proposed solution are MEAs that fully envelop the organoid [martinelli_e-flower_2024].
Another factor that necessitates deeper research is that of lacking reproducibility between organoids. Further methods such as those described by researchers [govindan_mass_2021] are essential for this field. Moreover, to enable longer testing and use of the organoids, research must find ways to minimize necrosis of brain organoid tissue. Possible paths to increased lifespans include improving the medium the organoids are immersed in. Alternatively, there have been experiments and proposals of fusing brain organoids with blood vessel organoids [koo_past_2019, sun_generation_2022, liu_unleashing_2024]. Further implementation of suchs approaches might lead to longer life times and thus more prolonged and complex studies involving more mature organoids. Such an advancement has the potential to increase the processing power of the brain organoids and enable more complex applications due to increased numbers of cells forming more complex structures [sun_generation_2022].
Further integration of AI tools in brain organoid experiments could help in characterizing and analyzing differences between organoids to cluster them according to behavior patterns and identify comparable groups [shi_organoid_2024]. More widescale research into this integration of brain organoid technology with traditional ANNs might also facilitate making use of deeper capabilities of brain organoids for computing.
All in all, research on classic neuromorphic systems and brain organoid computing should ideally work in tandem, with insights and discoveries of the function of BNNs guiding research and development of advanced neuromorphic systems. Though neuromorphic systems are based on brain functionality, research overlap between these deeply related fields remains limited as of today [basu_organoid_2024].
@@
Though researchers have compared models of consciousness and evaluated the sentience of brain organoids [montoya_what_2023, croxford_case_2024], no research regarding the opinions of wider stakeholders, such as the public, policymakers, or industry leaders has been conducted. Findings on this topic can influence future ethical conduct and ensure that further progress in this field takes public and regulatory opinion into account, minimizing the risk of adverse repercussions.
Platforms like FinalSpark’s Neuroplatform are designed to host and facilitate large scale computing experiments using brain organoids offering remote access to BNNs with associated MEAs and help future research in this area [jordan_open_2024]. This remotely accessible platform presents a novel opportunity for researchers worldwide to experiment with a standardized set of organoids and is already being used by several groups of researchers and universities. FinalSpark is also working on new applications of the technology, such as a demonstration of a brain organoid controlling a virtual butterfly in a 3D environment accessible to the public [burger_lab-grown_2024, finalspark_lab-grown_2024]. While not a holistic solution, this platform facilitates comprehensive studies on brain organoids, their intrinsic computing mechanisms, effects of adjusting environmental conditions, all the while collecting data that can help researchers with further efforts.
Further efforts are being conducted at Johns Hopkins University in its SURPASS program. This research is directed at several avenues of brain organoids and their computing capabilities, such as training organoids to play video games over the internet and control real-world robots [hartung_organoid_2024], with one long-term goal being the development of organoids consisting of up to one billion cells to better represent mature BNNs [johns_hopkins_university_first_2024].

### v2

@@ II-B4 Interfacing
Current brain organoid computing experiments mainly rely on observing changes in the pattern of elicited spikes [4]. Typically, predefined regions of activity are read out and mapped according to predetermined rule sets [7, 11]. Thus, the readout resolution is still quite limited and requires further research to make better use of brain organoids and their capabilities. The same holds true for applying input data to an organoid.
Most current experiments rely on reinforcing organoid activity with external electrical stimuli emulating rewards or punishments to train them to produce desired neural activity [7]. However, there is no general method of reliably performing input-output operations with brain organoids [4].
Current reinforcement training involves providing the brain organoids with predictable stimuli [^2] In this case 100Hz for 100ms. , typically with short, specific high-frequency signals, to targeted cell regions for positive feedback, and unpredictable stimuli [^3] In this case 5Hz for 4s., characterized by ”random” regions being stimulated with longer, low-frequency signals, for negative feedback[11].
In addition to this, random spikes elicited by neurons generate noise, which can result in unreliable and variable results [6], further contributing to the issue of interpreting neural activity and interfacing with the organoids. There is early interfacing software, which aims to solve these issues and provide the capabilities of performing described experiments. There is early software that aims to simplify interfacing with brain organoids, for example enabling the brain organoids to control external agents [43] and to be addressed remotely over the internet [4].
@@ II-B5 Ethics
The foremost ethical concern regarding experimentation on human brain organoids is the possibility of them being conscious. Teams of researchers have conducted analyses of experiments and research data to try to determine if current brain organoids display signs of higher-order thinking. Current research states that these organoids do not possess the complexity to allow for higher-order brain function, associated with consciousness [44]. Researchers also argue that the ability to learn and complex activity signatures by themselves is not a sign of consciousness and that without a corporeal state, consciousness cannot arise [45]. However, experiments showcasing brain organoids and their plasticity, characterize them as possessing a basic form of sentience [11]. This does not constitute consciousness, as the concept remains undefined with several competing theoretical models. Currently, the only widely acknowledged consciousness theory that allows such brain organoids to be classified as being some level of conscious, is the integrated information theory (IIT) [46]. It states that any system capable of integrating information and learning independently, is at least somewhat conscious [47].
The experiments mentioned before demonstrate this learning capability of brain organoids [11, 7, 43, 19]. Following the IIT, brain organoids qualify as being at least somewhat conscious [46]. However, this assessment itself is contested as some researchers in the field argue that the IIT and its representation is basically flawed for accurately depicting possible states of consciousness in in vitro cultures [48]. The definition of consciousness remains open, with numerous contradictory models and theories [47, 45], leaving the question unresolved whether brain organoids can feel pain or process complex emotions and sensations.
Due to ethical concerns, some countries, such as the U.S.A. and Germany, impose regulations on research with human brain organoids, which could limit research in the future. However, in said countries, no laws prohibit research with human brain organoids. Nonetheless, guidelines for experimentation with human stem cells, must be followed and monitored [49]. If consciousness is indeed discovered in more complex brain organoids, stricter regulations could slow or even halt progress and future organoid-based computers might be entirely prohibited.
@@ II-B6 Complexity
Though brain organoids exhibit complex neural structures, they are obviously by far not as complex as fully developed human brains and can be compared with early stages of embryonic development [39] at best. This limits their current processing potential. Thus, contemporary brain organoids are vastly inferior in terms of processing power to traditional processing methods and will most likely continue to be until more developed organoids can be grown.
In addition to this writing, reading, and interpreting information to and from such organoids remains an additional challenge that hinders current developments [19]. While certain promising capabilities of BOs have been demonstrated experimentally, the full integration and understanding of these systems remain major challenges. Thus, a better understanding of neural structures, their processing of stimuli, and the signals they generate is necessary. Current challenges of brain organoids for use in computing culminate in one word: complexity. The technology comes with various caveats, large challenges all by themselves, resulting in an incredibly complex set of problems to solve, if researchers want to fully understand and make use of brain organoids in a computing context, being the major stumbling block for current applications of brain organoids in a computational context.
@@ III-A Artificial Intelligence
Biological neural networks like human brains can be seen as larger, more complex and ANNs with an extremely high degree of interconnectivity [24], which formed the conceptual basis for digital systems such as ChatGPT, Google Gemini and others. The use of brain organoids, too, could heavily influence the field of AI in the next couple of years. In fact, human brain organoids have been used to conduct traditionally AI related tasks.
@@
Basic image processing [6] as well as the prediction of the behavior of dynamic systems exhibiting chaotic behavior [19] are typical tasks to which in vitro neurons have been applied.
@@
For the latter, such brain organoids required only 4 training epochs to achieve comparable predictive performance, whereas a machine learning algorithm with long short-term memory (LSTM) needed 50 epochs to outperform the biological system, representing a >>90% training time decrease [19].
@@
The demonstrated ability of human brain organoids to adapt to stimuli makes these functionally capable of solving more complex tasks typically associated with machine learning powered AI, such as speech recognition [19] or actively controlling systems such as video games [11].
@@
The inherent capability of organoids to change their connection structure makes them – at least in perspective – more powerful than traditional ANNs. Contemporary ANNs are larger and computationally intensive and therefore require pruning and specialization for efficient deployment [50]. In contrast to this, the self-organized, redundant, and interconnected nature of brain organoids lends itself to being better suited for generalist AI systems [24, 18]. Human brain organoids could therefore be trained in shorter periods, given some major advances in interfacing technology, provide efficient AI solutions, and complete tasks in parallel due to the neuronal structure that constitutes them. Their ability to learn autonomously opens possibilities for the use of this technology for a wide range of tasks that currently rely on machine learning, with BNNs potentially surpassing the abilities of contemporary artificial systems.
@@
Another area of application is the integration of classic ANNs with such organoids in order to complement each other’s strengths and weaknesses. This has been coined as organoid intelligence (OI) [3, 17].
@@
Analyzing vast amounts of data remains a strong suit of ANNs [51] and can be leveraged to characterize and analyze the vast amounts of data generated by brain organoids. This might be used in turn to better understand BNNs and their characteristics. Supplementing ANNs that excel at specialized tasks [18] with more generalized intelligences through the application of brain organoids might eventually lead to significant advances in the field of AI [19, 17]. Being able to learn from minimal sets of data, combined with the analytical strength of contemporary AI systems can enable superior predictive models, higher level decision-making and data analysis models [17] (organoid intelligence).
@@
Potential real world applications have been proposed and demonstrated in the field of robotics [52, 53, 54], AI algorithms in general [19, 3], such as those used for natural language processing, etc.
@@ III-B Neuromorphic Computing Research
Current neuromorphic processing units are limited by the available processing power with respect to the scope of the brain regions that they can currently simulate. Only partial (tiny) simulations of brain functionality [19] are possible today. Though brain organoids similarly do not represent fully developed brains and a fully diverse set of cells, current brain organoids are still more representative of portraying true brain functionality than any artificial means developed as of yet [39]. Accordingly, they can be used as inspiration for improved neuromorphic systems. The current and potential benefits of brain organoids can be leveraged for progress on more efficient, adaptable, and biologically representative neuromorphic processing units, see [55].
Contemporary research efforts in the area of neuromorphic processors almost exclusively stick to digital computer architectures, thus limiting the plasticity of the neuromorphic architectures under consideration [16]. However, there are several recent efforts to provide such systems with new substrates that can better mimic the plasticity of BNNs, with the use of memristors being one of the more widespread approaches [56, 16]. The ability of memristive devices to change their conductance according to basically the time integral over the current flowing through the device (there are devices exhibiting thresholds to distinguish between writing and reading) and store this state [56], makes them highly interesting for research on neuromorphic systems. Memristors are currently used in tasks requiring autonomous learning and it has been demonstrated that on-chip/on-the-edge learning is a viable area of application. One particular example showed memristor based integrated circuits which learned to control a remote controlled car with input provided through a camera [57]. Other examples include the recognition of written and spoken digits or words as well as the solution of dynamic systems [58, 56].
@@
Another key characteristic of biological systems is that of cell diversity. Contemporary neuromorphic architectures focus mainly on emulating neurons in a very simplified way, neglecting the computational functions provided by other cell types critical to cognition [16, 19]. For instance the actual behavior of dendrites is crucial in the processing of information within biological neuron, however, they are modeled in a simplified way, not fully emulating their actual function [59]. They have been shown to be paramount in enabling continued learning and long-term plasticity [59]. Employing better models could possibly alleviate the issue of catastrophic forgetting in ANNs, a problem for which currently no optimal solution exists [60].
In addition to this, quite a lot of different cell types are completely missing from current neuromorphic systems, such as astrocytes, which play important roles in network plasticity and regulating the spiking behavior of neurons [40]. Some experiments aim to simulate and integrate such missing cell types, though research in this direction is still in its infancy [40]. The plasticity of BNNs stems – at least partially – from these different cells and the inherent prediction of possible future stimuli [20, 17]. In contrast to these contemporary digital approaches, brain organoids already contain these cells, making them more biologically representative than contemporary neuromorphic systems [39, 11, 4].
Given the limitations of current neuromorphic systems in replicating biological complexities, their adaptability to uncertainty and real-world scenarios is significantly hampered. This is partly due to the absence of a comprehensive biological representation of neural structures [19]. Furthermore, while spiking associated with BNNs enhances their efficiency [16, 25], most of neuromorphic systems of today are not of the spiking type [58, 56, 57], limiting their potential efficiency, although there are exceptions [40, 16].
@@ IV Key Research
Maybe the most important area of research is the question of how to provide brain organoids with more abstract information and how to extract and interpret their responses. Solving these problems will advance their application in advanced AI use cases in general [4]. Consequently, research must focus on advancing MEA technology. This is crucial to more finely discern spikes from the organoids, in addition to allowing more precise addressing of their different cell structures. Current MEAs are severely limited with respect to the amount of addressable cells [19]. One proposed solution is MEAs that fully envelop the organoid [61, 62].
Another factor that necessitates deeper research is that of lacking reproducibility between organoids. Further methods such as those described by researchers [10] are essential for this field. Moreover, to enable longer testing and use of the organoids, research must find ways to minimize necrosis of brain organoid tissue. Possible paths to increased lifespans include improving the medium the organoids are immersed in. Alternatively, there have been experiments and proposals of fusing brain organoids with blood vessel organoids [2, 36, 26]. Further implementation of such approaches might lead to longer lifetimes and thus more prolonged and complex studies involving more mature organoids. Such an advancement has the potential to increase the processing power of the brain organoids and enable more complex applications due to increased numbers of cells forming more complex structures [36].
Further integration of AI tools in brain organoid experiments could help in characterizing and analyzing differences between organoids to cluster them according to behavior patterns and identify comparable groups [63]. Broader research into this integration of brain organoid technology with traditional ANNs might also facilitate making use of deeper capabilities of brain organoids for computing.
All in all, research on classic neuromorphic systems and brain organoid computing should ideally work in tandem, with insights and discoveries of the function of BNNs guiding research and development of advanced neuromorphic systems. Though neuromorphic systems are based on brain functionality, research overlap between these deeply related fields remains limited as of today [17].
@@
Though researchers have compared models of consciousness and evaluated the sentience of brain organoids [46, 45], no research regarding the opinions of wider stakeholders, such as the public, policymakers, or industry leaders has been conducted. Findings on this topic can influence future ethical conduct and ensure that further progress in this field takes public and regulatory opinion into account, minimizing the risk of adverse repercussions.
Platforms like FinalSpark’s Neuroplatform are designed to host and facilitate large-scale computing experiments using brain organoids offering remote access to BNNs with associated MEAs and help future research in this area [4]. This remotely accessible platform presents a novel opportunity for researchers worldwide to experiment with a standardized set of organoids and is already being used by several groups of researchers and universities. FinalSpark is also working on new applications of the technology, such as a demonstration of a brain organoid controlling a virtual butterfly in a 3D environment accessible to the public [64, 65]. While not a holistic solution, this platform facilitates comprehensive studies on brain organoids, their intrinsic computing mechanisms, effects of adjusting environmental conditions, all the while collecting data that can help researchers with further efforts.
Further efforts are being conducted at Johns Hopkins University in its SURPASS program. This research is directed at several avenues of brain organoids and their computing capabilities, such as training organoids to play video games over the internet and control real-world robots [54], with one long-term goal being the development of organoids consisting of up to one billion cells to better represent mature BNNs [9].